OpenStudy (anonymous):

The present population, P, in thousands, of a town is related to the initial population of the town at time t = 0 by the formulaP = 185(2)(0.05t)where t is measured in years. The town was populated in 1880. What was the population (to the nearest 1,000) in the year 2000?

OpenStudy (anonymous):

Is formula 185*2*0.05t

OpenStudy (anonymous):

?

OpenStudy (anonymous):

no im sorry the formula is 185(2)^(0.05t)

OpenStudy (anonymous):

I think you mean: P(t) = \[185(2)^{(0.05t)}\] so if you subtract 2000-1880 to get t then substitute that into the equation.

OpenStudy (anonymous):

t should be 120 120*.05 = 6 2^6 = 64 185*64 = 11840 people

OpenStudy (anonymous):

or 12,000 people if you round

OpenStudy (anonymous):

okay just to double check your answer, heres additional informatio. the population of the year 1880 was 185,000

OpenStudy (anonymous):

then in that case set 1880= 185(2)^(0.05t) and solve for t. Add 120 to the t value and substitute that value back into the P(t)

OpenStudy (anonymous):

and what did u come up with because im stuck

OpenStudy (dumbcow):

from 1880 to 2000, 120 years have passed so this will be the t_value you use to evaluate the population in year 2000 \[P = 185*2^{.05*120}\] \[P = 185*2^{6}\] \[P = 11,840\] since this is in thousands add 3 zeroes

OpenStudy (anonymous):

wouldnt you add 11,840 to the population 185,000 because thats how much the popultion increased over 120 years?

OpenStudy (dumbcow):

no because we left the 185 in the equation their initial population is already factored in the increase would be 11,840 - 185

OpenStudy (dumbcow):

in other words the equation gives population at any time t not just the increase

OpenStudy (anonymous):

damn. well then maybe idid it wrong and maybe the population of 1880 isnt 185,000 .....ughh im confusing myself.

OpenStudy (anonymous):

??????????????????

OpenStudy (dumbcow):

If what you posted above is correct then yes when t=0 in 1880 the population is 185,000